OpenStudy (anonymous):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = 18, b = 26, c = ? A. 31.16 B. 31.46 C. 31.62 D. 31.78

OpenStudy (johnweldon1993):

So if we have a right triangle with those sides  We can use the pythagorean theorem to solve for that hypotenuse... \[\large A^2 + B^2 = C^2\] Or more specifically \[\large C = \sqrt{A^2 + B^2}\] Just plug in your values to solve for 'c'

OpenStudy (anonymous):

is it c?

OpenStudy (girlgoyleh.):

for the people who'll need it, it is c